# A simple GPU implementation of spectral-element methods for solving 3D Poisson type equations on rectangular domains and its applications

### v2

@@ 1 Introduction
We remark that a similar simple implementation on GPUs can also be achieved using Python using the Python package JAX, which however provides better performance than MATLAB only for single precision computation. Our numerical tests and comparison on one Nvidia A100 GPU card suggest that MATLAB implementation performs better than Python for double precision computation for large problems like one billion DoFs.
@@ 2 A spectral-element method for Poisson type equations
with a constant coefficient $\alpha>0$ and homogeneous Neumann boundary conditions on a rectangular domain $\Omega$. We only consider the spectral-element method with continuous piecewise $Q^{k}$ polynomial basis on uniform rectangular meshes, and all integrals are approximated by $(k+1)$-point Gauss–Lobatto quadrature [14].
@@ 3.7 Comparison with implementation in Python
For implementing (14) on both CPU and GPU, similar to the implementation in MATLAB shown in Table 1, (14) can be implemented efficiently using the function jax.numpy.einsum in the Python package JAX as shown in Table 13.
@@
In Table 14, we list the online computational time comparison of similar implementations in MATLAB and Python on A100 for solving a 3D Poisson equation $200$ times. As we can see in Table 14, for double precision computation and problems with size smaller than $1000^{3}$, there is no significant difference in the online computational time between MATLAB and Python on GPUs. However, on A100 with 80G memory, MATLAB allows a problem size as large as $1250^{3}$, for which Python can handle only with single precision computation.
@@
Single precision Double precision Single precision Double precision
$200^{3}$ $4.80$E-$1$ $6.70$E-$1$ $4.72$E-$1$ $5.20$E-$1$
$250^{3}$ $5.85$E-$1$ $1.17$E00 $7.32$E-$1$ $9.11$E-$1$
$300^{3}$ $7.43$E-$1$ $2.10$E00 $1.54$E00 $1.80$E00
$350^{3}$ $1.39$E00 $3.47$E00 $2.56$E00 $3.04$E00
$400^{3}$ $1.51$E00 $5.47$E00 $4.63$E00 $5.07$E00
$450^{3}$ $2.89$E00 $8.74$E00 $7.05$E00 $8.09$E00
$500^{3}$ $2.92$E00 $1.20$E1111 $8.92$E00 $1.07$E1111
$550^{3}$ $6.07$E00 $1.82$E1111 $1.45$E1111 $1.65$E1111
$600^{3}$ $5.41$E00 $2.40$E1111 $1.97$E1111 $2.28$E1111
$650^{3}$ $1.02$E1111 $3.35$E1111 $3.02$E1111 $3.32$E1111
$700^{3}$ $9.76$E00 $4.24$E1111 $3.55$E1111 $4.04$E1111
$750^{3}$ $1.66$E1111 $5.66$E1111 $4.57$E1111 $5.22$E1111
$800^{3}$ $1.45$E1111 $7.14$E1111 $6.32$E1111 $6.89$E1111
$850^{3}$ $2.66$E1111 $9.28$E1111 $8.01$E1111 $9.05$E1111
$900^{3}$ $2.37$E1111 $1.15$E2222 $1.05$E2222 $1.15$E2222
$950^{3}$ $4.10$E1111 $1.37$E2222 $1.20$E2222 $1.35$E2222
$1000^{3}$ 3.173.173.173.$1$ $1.40$E2222 $1.60$E2222
$1050^{3}$ 6.126.126.126.$1$ $1.87$E2222 $2.04$E2222
$1100^{3}$ $4.77$E1111 - $2.16$E2222 $2.51$E2222
$1150^{3}$ $7.96$E1111 - $2.39$E2222 $2.91$E2222
$1200^{3}$ $6.30$E1111 - $2.95$E2222 $3.34$E2222
$1250^{3}$ $1.13$E2222 - $3.45$E2222 $4.13$E2222
Table 14: Online computational time on one Nvidia A100 80G GPU card of $Q^{5}$ SEM for solving a 3D Poisson equation $200$ times. The time unit is second. For double precision computation in Python on A100, an out-of-memory error will emerge for problems with size larger than $950^{3}$. MATLAB with single precision is not significantly faster than MATLAB with double precision, thus not shown.
As shown in Table 14, for larger problems such as one billion DoFs, the fastest implementation on GPU is Python with single precision computation, which might be suitable for some practical simulations. To see how accuracy is affected by single precision computation, we list some results in Table 15. In general, the implementation for high order SEM with single precision is not robust on A100, e.g., computation with SEM for the problem in Figure 8 might blow up.
$Q^{5}$ spectral-element method (single precision on Apple M1 CPU)
FEM Mesh Dirichlet boundary Neumann boundary
Total DoFs $\ell^{2}$ error order Total DoFs $\ell^{2}$ error order
$2^{3}$ $9^{3}$ $2.27$E-$1$ $11^{3}$ $4.76$E-$1$
$4^{3}$ $19^{3}$ $3.91$E-$3$ $5.86$ $21^{3}$ $5.49$E-$3$ $6.44$
$8^{3}$ $39^{3}$ $4.11$E-$5$ $6.57$ $41^{3}$ $4.32$E-$5$ $6.99$
$16^{3}$ $79^{3}$ $1.73$E-$6$ $4.57$ $81^{3}$ $1.54$E-$6$ $4.81$
$32^{3}$ $159^{3}$ $1.66$E-$6$ $0.05$ $161^{3}$ $2.07$E-$6$$0.42$
$Q^{5}$ spectral-element method (single precision with A100)
@@
Table 15: Accuracy tests under single precision in Python on Apple CPU (M1 chip) and Nvidia GPU A100 for the 3D Poisson equation (1) with $\alpha=1$. The actual accuracy of single precision computation depends very much on the hardware and version of hardware drivers. For Python, we implement the code under the environment JAX version 0.4.19 for both Apple CPU (M1 chip) and Nividia GPU (A100), with Driver Version 535.86.10 and CUDA Version 12.2 for A100. See Table 2 for the results of MATLAB with double precision on Nvidia GPU A100.
@@
Total DoFs Single precision Double precision
@@
Table 16: Accuracy tests for FFT implementation in Python on Nvidia GPU A100 for the 3D Poisson equation (1) with $u^{*}=\sin(2\pi x)\sin(3\pi y)\sin(4\pi z)$ and $\alpha=1$.
On the other hand, the second order finite difference implemented in Python Jax with single precision computation could be suitable for certain applications. For periodic boundary conditions, the eigenvectors of second order finite difference (i.e., $Q^{1}$ spectral-element method) can be implemented by FFT as shown in Table 17. As a demonstration, we include the computation results for the Cahn-Hilliard equation of Python in single precision on A100 in Figure 9, which is comparable to the double precision results on A100 in Figure 8.
@@
Figure 9: Snapshots of the same problem in Figure 8, implemented by Python in single precision on A100 for second order finite difference (i.e., $Q^{1}$ spectral-element method) by FFT with total DoFs $800^{3}$.
@@ 4 Concluding remarks
In this paper, we have discussed a simple MATLAB 2023 implementation for accelerating high order methods on GPUs. For large enough 3D problems, a speed-up of at least 60 can be achieved on Nvidia A100. In particular, solving a 3D Poisson type equation with one billion DoFs costs only 0.8 second for $Q^{k}$ spectral-element method. As examples of applications, We applied this fast solver to solve a linear (time-independent) Schrödinger equation and a nonlinear (time-dependent) Cahn-Hilliard equation in three-dimension.

### v3

@@ 1 Introduction
We remark that a similar simple implementation on GPUs can also be achieved using Python using the Python package JAX, which however provides better performance than MATLAB only for single precision computation under TensorFloat-32 precision format. Our numerical tests and comparison on one Nvidia A100 GPU card suggest that MATLAB implementation performs better than Python for double precision computation for large problems like one billion DoFs.
@@ 2 A spectral-element method for Poisson type equations
with a constant coefficient $\alpha>0$ and homogeneous Neumann boundary conditions on a rectangular domain $\Omega$. In this paper, we only consider the spectral-element method with continuous piecewise $Q^{k}$ polynomial basis on uniform rectangular meshes, and all integrals are approximated by $(k+1)$-point Gauss–Lobatto quadrature [14].
@@ 2.2 Inversion by eigenvalue decomposition
We remark that many other types of boundary conditions on a cubic domain can be written in the form of (3) or (5) include periodic boundaries and the simple homogeneous Dirichlet boundary treatment described in [14]. For non-homogeneous Neumann and Dirichlet boundary conditions, the boundary conditions will be added to the right hand side of (3) or (5), and the left hand side of (3) or (5) stays the same, i.e., the matrix to be inverted stays the same. The left hand side of (5) would no longer hold if the boundary value problem does not have a tensor product structure, e.g., on a curved domain.
@@ 3.7 Comparison with implementation in Python
For implementing (14) on both CPU and GPU, similar to the implementation in MATLAB shown in Table 1, (14) can be efficiently implemented using the function jax.numpy.einsum in the Python package JAX as shown in Table 13.
@@
In Table 14, we list the online computational time comparison of similar implementations in MATLAB and Python on A100 for solving a 3D Poisson equation $200$ times. As we can see in Table 14, for double precision computation and problems with size smaller than $1000^{3}$, there is no significant difference in the online computational time between MATLAB and Python on GPUs. However, on A100 with 80G memory, MATLAB allows a problem size as large as $1250^{3}$, for which Python can handle only with single precision computation. It is noteworthy that the performance of single precision computation in Python can be significantly affected by the use of the TF32 (TensorFloat-32) format on NVIDIA GPUs. TF32 is a lower-precision format that requires less memory bandwidth and compute resources compared to the traditional FP32 (Float-32) format, leading to faster computation at the cost of some loss in precision.
@@
Single(TF32) Single(FP32) Double Single Double
$200^{3}$ $4.80$E-$1$ $6.88$E-$1$ $6.70$E-$1$ $4.72$E-$1$ $5.20$E-$1$
$250^{3}$ $5.85$E-$1$ $9.92$E-$1$ $1.17$E00 $7.32$E-$1$ $9.11$E-$1$
$300^{3}$ $7.43$E-$1$ $1.67$E00 $2.10$E00 $1.54$E00 $1.80$E00
$350^{3}$ $1.39$E00 $2.80$E00 $3.47$E00 $2.56$E00 $3.04$E00
$400^{3}$ $1.51$E00 $4.56$E00 $5.47$E00 $4.63$E00 $5.07$E00
$450^{3}$ $2.89$E00 $7.24$E00 $8.74$E00 $7.05$E00 $8.09$E00
$500^{3}$ $2.92$E00 $9.48$E00 $1.20$E1111 $8.92$E00 $1.07$E1111
$550^{3}$ $6.07$E00 $1.41$E1111 $1.82$E1111 $1.45$E1111 $1.65$E1111
$600^{3}$ $5.41$E00 $1.99$E1111 $2.40$E1111 $1.97$E1111 $2.28$E1111
$650^{3}$ $1.02$E1111 $2.80$E1111 $3.35$E1111 $3.02$E1111 $3.32$E1111
$700^{3}$ $9.76$E00 $3.41$E1111 $4.24$E1111 $3.55$E1111 $4.04$E1111
$750^{3}$ $1.66$E1111 $4.48$E1111 $5.66$E1111 $4.57$E1111 $5.22$E1111
$800^{3}$ $1.45$E1111 $6.12$E1111 $7.14$E1111 $6.32$E1111 $6.89$E1111
$850^{3}$ $2.66$E1111 $7.92$E1111 $9.28$E1111 $8.01$E1111 $9.05$E1111
$900^{3}$ $2.37$E1111 $1.00$E2222 $1.15$E2222 $1.05$E2222 $1.15$E2222
$950^{3}$ $4.10$E1111 $1.17$E2222 $1.37$E2222 $1.20$E2222 $1.35$E2222
$1000^{3}$ $3.17$E1111 $1.40$E2222 - $1.40$E2222 $1.60$E2222
$1050^{3}$ $6.12$E1111 $1.81$E2222 - $1.87$E2222 $2.04$E2222
$1100^{3}$ $4.77$E1111 $2$ $2.16$E2222 $2.51$E2222
$1150^{3}$ $7.96$E1111 $2$ $2.39$E2222 $2.91$E2222
$1200^{3}$ $6.30$E1111 $2$ $2.95$E2222 $3.34$E2222
$1250^{3}$ $1.13$E2222 $3.36$E2222 - $3.45$E2222 $4.13$E2222
Table 14: Online computational time of single precision and double precison on one Nvidia A100 80G GPU card, for $Q^{5}$ SEM for solving a 3D Poisson equation $200$ times. The time unit is second. For double precision computation in Python on A100, an out-of-memory error will emerge for problems with size larger than $950^{3}$.
As shown in Table 14, for larger problems such as one billion DoFs, the fastest implementation on GPU is Python with single precision computation, which might be suitable for some practical simulations. When using the default TF32 setting on A100, the accuracy of the $Q^{5}$ spectral-element method deteriorates for larger problem sizes, with the order of accuracy dropping below the expected value. However, by setting the environment variable NVIDIA_TF32_OVERRIDE=0 to disable TF32 precision for certain operations, such as matrix multiplication and jax.numpy.einsum, the accuracy can be improved, and the expected order of accuracy is maintained. To investigate the impact of TF32 on the accuracy of single precision computation in Python, we conducted additional tests with and without the TF32 format. Table 15 presents the accuracy results for the $Q^{5}$ spectral-element method under single precision on A100 with both TF32 and FP32 format. In general, the implementation for high order SEM with TF32 single precision is not robust on A100, e.g., computation with SEM for the problem in Figure 8 might blow up.
$Q^{5}$ spectral-element method (TF32 single precision)
@@
$Q^{5}$ spectral-element method (FP32 single precision)
FEM Mesh Dirichlet boundary Neumann boundary
Total DoFs $\ell^{2}$ error order Total DoFs $\ell^{2}$ error order
$2^{3}$ $9^{3}$ $2.27$E-$1$ $11^{3}$ $4.76$E-$1$
$4^{3}$ $19^{3}$ $3.91$E-$3$ $5.86$ $21^{3}$ $5.49$E-$3$ $6.44$
$8^{3}$ $39^{3}$ $4.11$E-$5$ $6.57$ $41^{3}$ $4.32$E-$5$ $6.99$
$16^{3}$ $79^{3}$ $1.67$E-$6$ $4.62$ $81^{3}$ $1.63$E-$6$ $4.72$
$32^{3}$ $159^{3}$ $1.34$E-$6$ $0.32$ $161^{3}$ $1.95$E-$6$$0.26$
Table 15: Accuracy tests under TF32 and FP32 single precision in Python on Nvidia GPU A100 for the 3D Poisson equation (1) with $\alpha=1$. The actual accuracy of single precision computation depends very much on the hardware and version of hardware drivers. For Python, we implement the code under the environment JAX version 0.4.19 for Nvidia GPU A100, with Driver Version 535.86.10 and CUDA Version 12.2. See Table 2 for the results of MATLAB with double precision on Nvidia GPU A100.
@@
Total DoFs TF32 Single precision Double precision
@@
Table 16: Accuracy tests for second order finite difference with periodic boundary (FFT implementation) in Python on Nvidia GPU A100 for the 3D Poisson equation (1) with $u^{*}=\sin(2\pi x)\sin(3\pi y)\sin(4\pi z)$ and $\alpha=1$.
On the other hand, the second order finite difference implemented in Python Jax with TF32 single precision computation is robust as suggested by Table 16. For periodic boundary conditions, the eigenvectors of second order finite difference (i.e., $Q^{1}$ spectral-element method) can be implemented by FFT as shown in Table 17. As a demonstration, we include the computation results for the Cahn-Hilliard equation of Python in TF32 single precision on A100 in Figure 9, which is comparable to the double precision results on A100 in Figure 8.
@@
Figure 9: Snapshots of the same problem in Figure 8, implemented by Python in TF32 single precision on A100 for second order finite difference (i.e., $Q^{1}$ spectral-element method) by FFT with total DoFs $800^{3}$.
@@ 4 Concluding remarks
In this paper, we have discussed a simple MATLAB 2023 implementation for accelerating high order methods on GPUs. For large enough 3D problems, a speed-up of at least 60 can be achieved on Nvidia A100. In particular, solving a 3D Poisson type equation with one billion DoFs costs only 0.8 second for $Q^{k}$ spectral-element method. As examples of applications, we applied this fast solver to solve a linear (time-independent) Schrödinger equation and a nonlinear (time-dependent) Cahn-Hilliard equation in three-dimension. We expect the proposed simple implementation to have the same performance for any problem with similar tensor product structure, e.g., exponential time differencing and spectral fractional Laplacians.